# Quadratic-form Optimal Transport

### v2

@@ Abstract
We introduce the framework of quadratic-form optimal transport (QOT), whose transport cost has the form $\iint c\,\mathrm{d}\pi\otimes\mathrm{d}\pi$ for some coupling $\pi$ between two marginals. Interesting examples of quadratic-form transport cost and their optimization include inequality measurement, the variance of a bivariate function, covariance, Kendall’s tau, the Gromov–Wasserstein distance, quadratic assignment problems, and quadratic regularization of classic optimal transport. QOT leads to substantially different mathematical structures compared to classic transport problems and many technical challenges. We illustrate the fundamental properties of QOT, provide several cases where explicit solutions are obtained, and give general lower bounds of the optimal transport costs. For a wide class of cost functions, including the rectangular cost functions, the QOT problem is solved by a new coupling called the diamond transport, whose copula is supported on a diamond in the unit square.
@@ 1 Introduction
to minimize $\mbox{to minimize\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ }\mathbb{E}[(\theta_{1}|X-X^{\prime}|+\theta_{2}|Y-Y^{\prime}|)^{2}]\mbox{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ over $\pi\in\Pi(\mu,\nu)$},$ (1.1)
where $(X,Y)$ and $(X^{\prime},Y^{\prime})$ are independently drawn from the distribution $\pi$ [^1] This problem has a Kantorovich formulation. For a practical application with finitely many individuals, one may further require $\pi$ to be induced by a permutation (that is, the Monge formulation), but we will mainly focus on the Kantorovich formulation in our study, which is technically tractable and offers approximations for optimizers in the Monge setting (see Section 5).
@@
To analyze (1.1), one may first look at some commonly encountered couplings. A quick observation is that if $\mu=\nu$ and $\pi$ is the comonotone coupling $\pi_{\rm com}$ (formal definition in Section 2), then $Y=X$ and $Y^{\prime}=X^{\prime}$ (almost surely), and thus $|X-X^{\prime}|=|Y-Y^{\prime}|$. Since the distributions of $|X-X^{\prime}|$ and $|Y-Y^{\prime}|$ are determined solely by $\mu$ and $\nu$, the transport cost (the expectation in (1.1)) is maximized by $\pi_{\rm com}$ in this case. From there, one may then wonder whether the antimonotone coupling $\pi_{\rm ant}$ minimizes (1.1). However, by choosing $\mu=\nu$ as the standard uniform distribution, $\pi_{\rm ant}$ also yields $|X-X^{\prime}|=|Y-Y^{\prime}|$, thus also maximizing transport cost. Therefore, intuitively, the optimal coupling must lie somewhere between the most positive coupling $\pi_{\rm com}$ and the most negative coupling $\pi_{\rm ant}$. It turns out that the optimal coupling, solved in full generality in Section 5, has a special structure that is different from both independence and mixtures of $\pi_{\rm com}$ and $\pi_{\rm ant}$.
@@
$c(x,y,x^{\prime},y^{\prime})=h(f(x,x^{\prime}),g(y,y^{\prime}))\mbox{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ for some real-valued bivariate functions $f,g,h$},$ (1.3)
@@
$c(x,y,x^{\prime},y^{\prime})=h(f(x,y),g(x^{\prime},y^{\prime}))\mbox{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ for some real-valued bivariate functions $f,g,h$}.$ (1.4)
@@
$c(x,y,x^{\prime},y^{\prime})=f(x,y)+g(x^{\prime},y^{\prime})\mbox{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ for some bivariate functions $f,g$};$
@@
$c(x,y,x^{\prime},y^{\prime})=\frac{1}{2}(f(x,y)-f(x^{\prime},y^{\prime}))^{2}\mbox{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ for some bivariate function $f$};$
@@
$c(x,y,x^{\prime},y^{\prime})=f(x,x^{\prime})g(y,y^{\prime})\mbox{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ for some bivariate functions $f,g$}.$
@@
where $\{a_{ij}\}_{1\leqslant i,j\leqslant n},\{b_{ij}\}_{1\leqslant i,j\leqslant n}$ are given $n\times n$ matrices and $S_{n}$ is the set of all permutations of $[n]=\{1,\dots,n\}$. This includes many prominent examples such as the traveling salesman problem. Our theory of QOT in the discrete case includes but is much more general than QAP. For instance, if the Monge assumption is relaxed, the optimizer may not even be a (deterministic) assignment; see Example 1.1 below.
@@
Figure 1: Illustration of the support of the comonotone transport $\pi_{\rm com}$, the antimonotone transport $\pi_{\rm ant}$, the X-transport $\pi_{\rm x}^{\lambda}$ with $\lambda\in(0,1)$, the V-transport $\pi_{\rm v}$, and the diamond transport $\pi_{\rm dia}$, which appear as QOT minimizers with marginals normalized to uniform distributions on $[0,1]$; for their corresponding cost functions and marginals, see Table 1 below. Blue lines indicate the support of the transport plans. The transport plans in (c), (d), and (e) appear new compared to classic OT. Precise definitions are given in Section 2 and Definitions 4.9 and 5.1.
@@
$(f(x,y)+a_{1})(f(x^{\prime},y^{\prime})+a_{2})$ $f$ submodular, $a_{1},a_{2}\in\mathbb{R}$ general $\pi_{\rm x}^{\lambda}=\lambda\pi_{\rm com}+(1-\lambda)\pi_{\rm ant}$ for some $\lambda\in[0,1]$ Proposition 4.3
@@
[…]
$f(|x-x^{\prime}|)g(y,y^{\prime})$ $f$ nonnegative increasing $g$ increasing supermodular and some regularity conditions $\mu$ is uniform on an interval $\pi_{\rm v}$ Theorem 4.10
@@
Table 1: A selected list of explicitly solved examples of QOT problems on the real line. The minimizers may not be unique. Certain moment assumptions on the marginals are omitted (compactness of support is sufficient). For definitions of the couplings, see Section 2 and Definitions 4.9 and 5.1. In Examples 4.56.8, many more explicit cost functions, some of which belong to the above general classes, are presented. The conclusions remain the same if QOT-irrelevant terms like $w_{1}(x,x^{\prime})+w_{2}(y,y^{\prime})+w_{3}(x,y^{\prime})+w_{4}(x^{\prime},y)$ are added to the cost function $c$ (Fact 2.1).
In Section 2, we formally present the framework of QOT on general Polish spaces. Section 3 provides general results on QOT, including properties of the optimizers and general lower bounds of the QOT cost. Section 4 examines several explicit solvable cases. We show that the comonotone coupling $\pi_{\rm com}$ and the antimonotone coupling $\pi_{\rm ant}$, as well as their mixtures, form solutions to many classes of cost functions, including quadratic costs, jointly submodular/supermodular costs, and Gromov–Wasserstein-type costs.
@@
In Section 6, we discuss many relevant examples of QOT in optimization, economics, computational OT, and statistics. Section 7 concludes with a few open questions.
The appendices contain further discussions, additional results, and omitted proofs. Appendix A gives a quadratic programming formulation of QOT. Appendix B shows that the independent coupling is rarely, but possibly, an optimizer of the QOT, and gives some interesting examples. In Appendix C, we consider the class of linear-exponential distance cost functions of the form $c(x,y,x^{\prime},y^{\prime})=|y-y^{\prime}|e^{-\gamma|x-x^{\prime}|},\leavevmode\nobreak\ \gamma>0$. This class of cost functions are minimized by the comonotone coupling, but its maximizers have interesting limiting behavior as $\gamma$ goes to 00 or $\infty$, such as the diamond and independent couplings, in some special senses. In particular, the limitting case $\gamma\to\infty$ is connected to recently studied measures of association (Chatterjee (2021); Deb et al. (2020)). Appendix D contains a more detailed discussion of the open questions from Section 7. Appendices EG collect omitted proofs of the results from Sections 25.
@@ 3 General properties
The results presented in this section hold for probability measures on general Polish spaces, except for the stability of QOT (Proposition 3.5), where we require $\mathfrak{X}=\mathfrak{Y}=\mathbb{R}$.
Convexity is an essential issue in classic OT theory, giving rise to many useful techniques such as duality and $c$-cyclical monotonicity. We start with a simple result that gives a sufficient condition for QOT to be convex. For a non-empty set $S$, we say that a symmetric function $\phi:S\times S\to\mathbb{R}$ is positive definite if for any $n\in\mathbb{N},\leavevmode\nobreak\ s_{1},\dots,s_{n}\in S$, and $c_{1},\dots,c_{n}\in\mathbb{R}$, it holds
@@
which is a generalization of positive semi-definite matrices.
@@
[…]
Despite the above result, the majority of QOT problems are not convex in $\pi$. This non-convex structure of QOT prohibits the use of classic tools such as duality. [^5] Notable exceptions include Vayer (2020, Theorem 4.2.5) and (Zhang et al., 2024, Theorem 1), and the latter result also analyzes sample complexity of the $(2,2)$-GW distance; see Example 6.6 below. In the rest of this section, we study the fundamental properties of QOT, which may not be convex. Specifically, we show that under certain assumptions, minimizers exist, the Monge optimal transport cost is equivalent to the Kantorovich one, QOT on $\mathbb{R}$ is stable, and the independent coupling is rarely an optimizer.
@@
Fact 3.3.
If $c\in\mathcal{C}(\mu,\nu)$, then the infimum of the quadratic-form transport cost in (2.1) is well-defined and not $\infty$- ∞.
The next result gives conditions under which minimizers of QOT exist.
@@
$\min_{\pi\in\Pi(\mu,\nu)}\iint c\,\mathrm{d}\pi\otimes\mathrm{d}\pi=\inf_{\pi\in\mathcal{T}(\mu,\nu)}\iint c\,\mathrm{d}\pi\otimes\mathrm{d}\pi.$
@@
We next show that similar to classic OT, QOT satisfies stability with respect to the marginals.
@@
Stability is crucial in classic OT theory as it ensures that numerical algorithms for solving the OT problem converge consistently. On the other hand, numerically computing or approximating the QOT solution remains a difficult task, as the discretized version remains an NP-hard problem (Loiola et al. (2007)). As a discrete Monge version of QOT, QAP provides feasible heuristic algorithms; see Burkard et al. (2012) and Çela (2013). However, one needs to be careful here since discrete QOT may not always have a Monge minimizer even if a bijective transport map exists (see Example 1.1). We leave the computational aspects of QOT for further investigation.
@@
The lower bounds in Propositions 3.6 and 3.7 are generally not sharp, but they are useful in proving the optimality of some transport plans.

### v4

@@ Abstract
We introduce the framework of quadratic-form optimal transport (QOT), whose transport cost has the form $\iint c\,\mathrm{d}\pi\otimes\mathrm{d}\pi$ for some coupling $\pi$ between two marginals. Interesting examples of quadratic-form transport cost and their optimization include inequality measurement, the variance of a bivariate function, covariance, Kendall’s tau, the Gromov–Wasserstein distance, quadratic assignment problems, and quadratic regularization of classic optimal transport. QOT leads to substantially different mathematical structures compared to classic transport problems and many technical challenges. We illustrate the fundamental properties of QOT and provide several cases where explicit solutions are obtained. For a wide class of cost functions, including the rectangular cost functions, the QOT problem is solved by a new coupling called the diamond transport, whose copula is supported on a diamond in the unit square.
@@ 1 Introduction
$\mbox{to minimize\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ }\mathbb{E}[(\theta_{1}|X-X^{\prime}|+\theta_{2}|Y-Y^{\prime}|)^{2}]\mbox{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ over $\pi\in\Pi(\mu,\nu)$},$ (1.1)
where $(X,Y)$ and $(X^{\prime},Y^{\prime})$ are independently drawn from the distribution $\pi$. [^1] This problem has a Kantorovich formulation. For a practical application with finitely many individuals, one may further require $\pi$ to be induced by a permutation (that is, the Monge formulation), but we will mainly focus on the Kantorovich formulation in our study, which is technically tractable and offers approximations for optimizers in the Monge setting (see Section 6).
@@
To analyze (1.1), one may first look at some commonly encountered couplings. A quick observation is that if $\mu=\nu$ and $\pi$ is the comonotone coupling $\pi_{\rm com}$ (formal definition in Section 2), then $Y=X$ and $Y^{\prime}=X^{\prime}$ (almost surely), and thus $|X-X^{\prime}|=|Y-Y^{\prime}|$. Expanding (1.1), we see that the transport cost (the expectation in (1.1)) consists of terms that do not depend on $\pi$, plus $2\theta_{1}\theta_{2}\mathrm{Cov}(|X-X^{\prime}|,|Y-Y^{\prime}|)$. Since the distributions of $|X-X^{\prime}|$ and $|Y-Y^{\prime}|$ are determined solely by $\mu$ and $\nu$, the covariance and hence the transport cost is maximized when $|X-X^{\prime}|=|Y-Y^{\prime}|$. From there, one may then wonder whether the antimonotone coupling $\pi_{\rm ant}$ minimizes (1.1). However, by choosing $\mu=\nu$ as the standard uniform distribution, $\pi_{\rm ant}$ also yields $|X-X^{\prime}|=|Y-Y^{\prime}|$, thus also maximizing transport cost. Therefore, intuitively, the optimal coupling must lie somewhere between the most positive coupling $\pi_{\rm com}$ and the most negative coupling $\pi_{\rm ant}$. It turns out that the optimal coupling, solved in full generality in Section 6, has a special structure that is different from both independence and mixtures of $\pi_{\rm com}$ and $\pi_{\rm ant}$.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
where $\{a_{ij}\}_{1\leqslant i,j\leqslant n},\{b_{ij}\}_{1\leqslant i,j\leqslant n}$ are given $n\times n$ matrices and $S_{n}$ is the set of all permutations of $n]=\{1,\dots,n\}$. This includes many prominent examples such as the traveling salesman problem. Our theory of QOT in the discrete case includes but is much more general than QAP. For instance, if the Monge assumption is relaxed, the optimizer may not even be a (deterministic) assignment; see Example 1.1 below.
@@
Figure 1: Illustration of the support of the comonotone transport $\pi_{\rm com}$, the antimonotone transport $\pi_{\rm ant}$, the X-transport $\pi_{\rm x}^{\lambda}$ with $\lambda\in(0,1)$, the V-transport $\pi_{\rm v}$, and the diamond transport $\pi_{\rm dia}$, which appear as QOT minimizers with marginals normalized to uniform distributions on $[0,1]$; for their corresponding cost functions and marginals, see Table 1 below. Blue lines indicate the support of the transport plans. The transport plans in (c), (d), and (e) appear new compared to classic OT. Precise definitions are given in Section 2 and Definitions 5.9 and 6.1.
@@
$(\alpha_{1}f(x,y)+C_{1})(\alpha_{2}f(x^{\prime},y^{\prime})+C_{2})$ $f$ submodular, $C_{1},C_{2},\alpha_{1},\alpha_{2}\in\mathbb{R}$ general $\pi_{\rm x}^{\lambda}=\lambda\pi_{\rm com}+(1-\lambda)\pi_{\rm ant}$ for some $\lambda\in[0,1]$ Proposition 5.3
@@
$(|x-x^{\prime}|,|y-y^{\prime}|) hh$ submodular $\nu=\mu\circ\ell^{-1}$ $\ell$ linear map $\pi_{\rm com}$ ($\ell$ increasing) $\pi_{\rm ant}$ ($\ell$ decreasing) Theorem 5.6
$(|x-x^{\prime}|)g(y,y^{\prime}) ff$ nonnegative increasing $g$ increasing supermodular and some regularity conditions $\mu$ is uniform on an interval $\pi_{\rm v}$ Theorem 5.10
@@
Table 1: A selected list of explicitly solved examples of QOT problems on the real line. The minimizers may not be unique. Certain moment assumptions on the marginals are omitted (compactness of support is sufficient). For definitions of the couplings, see Section 2 and Definitions 5.9 and 6.1. In Examples 3.15.5, many more explicit cost functions, some of which belong to the above general classes, are presented. The conclusions remain the same if QOT-irrelevant terms like $w_{1}(x,x^{\prime})+w_{2}(y,y^{\prime})+w_{3}(x,y^{\prime})+w_{4}(x^{\prime},y)$ are added to the cost function $c$ (Fact 2.1).
In Section 2, we formally present the framework of QOT on general Polish spaces. Section 3 collects many relevant examples of QOT in optimization, economics, computational OT, and statistics.
Section 4 provides general results on QOT, including properties of the optimizers and general lower bounds of the QOT cost, some of which extend known results on the GW distance. Section 5 examines several explicit solvable cases. We show that the comonotone coupling $\pi_{\rm com}$ and the antimonotone coupling $\pi_{\rm ant}$, as well as their mixtures, form solutions to many classes of cost functions, including quadratic costs, jointly submodular/supermodular costs, and Gromov–Wasserstein-type costs.
@@
Section 7 concludes with a few open questions. The appendices contain further discussions, additional results, and omitted proofs. Appendix A gives a quadratic programming formulation of QOT. Appendix B shows that the independent coupling is rarely, but possibly, an optimizer of the QOT, and gives some interesting examples. In Appendix C, we consider the class of linear-exponential distance cost functions of the form $c(x,y,x^{\prime},y^{\prime})=|y-y^{\prime}|e^{-\gamma|x-x^{\prime}|},\leavevmode\nobreak\ \gamma>0$. This class of cost functions are minimized by the comonotone coupling, but its maximizers have interesting limiting behavior as $\gamma$ goes to 0 or $\infty$, such as the diamond and independent couplings, in some special senses. In particular, the limitting case $\gamma\to\infty$ is connected to recently studied measures of association (Chatterjee (2021); Deb et al. (2020)). Appendix D contains a more detailed discussion of the open questions from Section 7. Appendices EG collect omitted proofs of the results from Sections 46.
@@ 2 Framework
3 Examples of QOT
In this section, we discuss several examples of QOT that appear in different fields and how they connect to the results in this work.
Example 3.1 (Sum of bivariate functions).
Example 3.2 (Variance minimization with given marginals).
$\begin{aligned} \iint c\,\mathrm{d}\pi\otimes\mathrm{d}\pi&=\begin{cases}\mathbb{E}[f(X,Y)^{2}]-\mathbb{E}[f(X,Y)]^{2}&\text{ if }\mathbb{E}[f(X,Y)^{2}]<\infty\\ \infty&\mbox{ otherwise}\end{cases}\\ &=\mathrm{Var}(f(X,Y)).\end{aligned}$ (3.1)
Example 3.3 (Covariance).
For $(X,Y)\buildrel\mathrm{law}\over{\sim}\pi$, we can verify
$=\mathbb{E}[XY]-\mathbb{E}[X]\mathbb{E}[Y]=\mathrm{Cov}(X,Y).$
Example 3.4 (Kendall’s tau).
$\tau=\mathbb{E}[\mathrm{sgn}((X-X^{\prime})(Y-Y^{\prime}))],$
where $(X^{\prime},Y^{\prime})$ is an independent copy of $(X,Y)$. Intuitively, it equals the probability of concordance minus that of discordance between $(X,Y)$ and $(X^{\prime},Y^{\prime})$. Clearly, $\tau$ of $(X,Y)\buildrel\mathrm{law}\over{\sim}\pi$ can be written as the quadratic-form transport cost with the type-XX cost function $c(x,y,x^{\prime},y^{\prime})=\mathrm{sgn}(x-x^{\prime})\,\mathrm{sgn}(y-y^{\prime})$. For given marginals $\mu,\nu$, it is well-known that $\tau(\pi)$ over $\pi\in\Pi(\mu,\nu)$ is maximized by the comonotone coupling with maximum value $1$ and minimized by the antimonotone coupling with minimum value 1-1 (this can also be checked by Theorem 5.4; see Example 5.5). Another equivalent formulation is $\tau(\pi)=4\int\pi\mathrm{d}\pi-1,$ from which the quadratic form in $\pi$ is visible.
Example 3.5 (Gini deviation and Gini coefficient).
where $Z^{\prime}$ is an independent copy of $Z$. The value $\mathrm{GD}(Z)$ is called the Gini deviation of $Z$, and $\mathrm{GC}(Z)$ is called the Gini coefficient of $Z$, both of which are commonly used as measures of distributional variability or inequality in economics and risk management; see e.g., Gastwirth (1971) and Furman et al. (2017). Similarly to the variance in (3.1), minimization of $\mathrm{GD}(f(X,Y))$ for some measurable $f:\mathbb{R}^{2}\to\mathbb{R}$ over $X\buildrel\mathrm{law}\over{\sim}\mu$ and $Y\buildrel\mathrm{law}\over{\sim}\nu$ can be written as the QOT problem with the type-XY cost function $|f(x,y)-f(x^{\prime},y^{\prime})|$. Moreover, the minimization of $\mathrm{GC}(X+Y)$ can be written as
which is equivalent to a QOT problem with cost function $c(x,y,x^{\prime},y^{\prime})=|x+y-x^{\prime}-y^{\prime}|$, noting that the denominator of (3.2) does not involve $\pi$. A minimizer of (3.2) is given by the antimonotone coupling $\pi_{\mathrm{ant}}$; see Example 5.5(v) below. This cost function is both type-XX and type-XY.
Example 3.6 (Gromov–Wasserstein (GW) distance).
A special case of QOT is the GW distance, a measure of the distance (or similarity) between two metric measure spaces introduced and studied by Mémoli (2007, 2011a). Suppose that $(\mathfrak{X},d_{\mathfrak{X}},\mu)$ and $(\mathfrak{Y},d_{\mathfrak{Y}},\nu)$ are metric measure spaces. For $p,q\geqslant 1$, the $(p,q)$-GW distance is defined as
This is an increasing transform of the minimum transport cost of QOT with a type-XX cost function. The GW distance satisfies the triangle inequality and defines a pseudo-metric on metric measure spaces (and a metric on isomorphism classes of metric measure spaces; see Sturm (2023, Corollary 9.3)). The GW distance is a widely used technique in data science, machine learning, computer vision, and computer graphics to align heterogeneous data sets or images (Mémoli (2011a); Peyré and Cuturi (2019)). However, in general, solving for the GW distance is a challenging task. In Theorems 6.2 and 6.9, we explicitly characterize the maximizers of the transport cost that appears in (3.3) for $p=2$, $q\in[1,2]$, $\mathfrak{X}=\mathfrak{Y}=\mathbb{R}$ equipped with the Euclidean distance, and symmetric marginals, where the symmetry is not needed for $q=1$. The GW literature also incorporates some earlier ideas discussed in this paper, such as existence of minimizers and Monge–Kantorovich equivalence (Proposition 4.4), lower bounds on the transport cost (Propositions 4.6 and 4.7), and connections to QAP. These ideas also manifest in studies of extensions of the GW distance (Arya et al. (2024); Bauer et al. (2024); Chowdhury and Mémoli (2019); Mémoli (2011b, 2012); Mémoli et al. (2023)), sometimes leading to closed-form solutions.
Example 3.7 (Quadratic Assignment Problem (QAP)).
If the probability measures $\mu,\nu$ are each uniformly distributed on $N$ points, the Monge QOT problem reduces to QAP, which was first introduced by Koopmans and Beckmann (1957) as a model for the allocation problem of indivisible economic activities. This connection is well known in the GW literature; see, for instance, Remark 4.6 of Mémoli (2011a). Recall the Koopmans–Beckmann problem from (1.5). The work of Lawler (1963) proposed a generalization of the Koopmans–Beckmann QAP, where one solves
$\min_{\sigma\in S_{n}}\sum_{i=1}^{n}\sum_{j=1}^{n}d_{ij\sigma_{i}\sigma_{j}},$
subject to $\pi\in\Pi(\mu,\nu),$
@@ 4 General properties
We now proceed to study general properties of QOT. The results presented in this section hold for probability measures on general Polish spaces, except for the stability of QOT (Proposition 4.5), where we require $\mathfrak{X}=\mathfrak{Y}=\mathbb{R}$.
Convexity is an essential issue in classic OT theory, giving rise to many useful techniques such as duality and $c$-cyclical monotonicity. We start with a simple result that gives a sufficient condition for QOT to be convex. For a non-empty set $S$, we say that a symmetric kernel $\phi:S\times S\to\mathbb{R}$ is a positive definite kernel if for any $n\in\mathbb{N},\leavevmode\nobreak\ s_{1},\dots,s_{n}\in S$, and $c_{1},\dots,c_{n}\in\mathbb{R}$, it holds
@@
which is a generalization of positive semi-definite matrices. A closely related concept is that of positive definite functions, that is, functions $\varphi:\mathbb{R}\to\mathbb{R}$ satisfying that $(x,y)\mapsto\varphi(x-y)$ is a positive definite kernel on $\mathbb{R}^{2}$.
@@
Denote by $\|\cdot\|$ the Euclidean norm. For ${\mathbf{z}},{\mathbf{z}}^{\prime}\in\mathbb{R}^{d}$, the kernels $({\mathbf{z}},{\mathbf{z}}^{\prime})\mapsto e^{-\alpha\|{\mathbf{z}}-{\mathbf{z}}^{\prime}\|}$ and $({\mathbf{z}},{\mathbf{z}}^{\prime})\mapsto e^{-\alpha\|{\mathbf{z}}-{\mathbf{z}}^{\prime}\|^{2}}$ are both positive definite for $\alpha>0$. Therefore, the QOT problem with the type-XX cost function $e^{-\alpha\|(x,y)-(x^{\prime},y^{\prime})\|}$ or $e^{-\alpha\|(x,y)-(x^{\prime},y^{\prime})\|^{2}}$ is convex. We will see more examples in Example 6.7, where we also show that QOT with cost function $e^{-\alpha\|(x,y)-(x^{\prime},y^{\prime})\|^{2}},\,\alpha\in(0,1/2]$ is minimized by the diamond transport.
Despite the above result, the majority of QOT problems are not convex in $\pi$. This non-convex structure of QOT prohibits the use of classic tools such as duality. [^5] Notable exceptions include Vayer (2020, Theorem 4.2.5) and (Zhang et al., 2024, Theorem 1), and the latter result also analyzes sample complexity of the $(2,2)$-GW distance; see Example 3.6. In the rest of this section, we study the fundamental properties of QOT, which may not be convex. Specifically, we show that under certain assumptions, minimizers exist, the Monge optimal transport cost is equivalent to the Kantorovich one, QOT on $\mathbb{R}$ is stable, and the independent coupling is rarely an optimizer.
@@
Fact 4.3.
If $c\in\mathcal{C}(\mu,\nu)$, then the infimum of the quadratic-form transport cost in (2.1) is well-defined and not -$\infty$.
The next result gives conditions under which minimizers of QOT exist and under which Monge is equivalent to Kantorovich. Similar results have been established for special cases such as the GW distance (Mémoli (2011a, Corollary 10.1) and Mémoli and Needham (2024, Theorem 3.2); see Example 3.6 for the formulation) and its extensions (e.g., Bauer et al. (2024, Theorem 2)). We also refer to Section 3 of Mémoli and Needham (2024) for further results that compare the Monge and Kantorovich problems in the case of GW costs.
@@
$\min_{\pi\in\Pi(\mu,\nu)}\iint c\,\mathrm{d}\pi\otimes\mathrm{d}\pi={\inf_{T\in\mathcal{T}(\mu,\nu)}\iint c(x,T(x),x^{\prime},T(x^{\prime}))\,\mathrm{d}\mu(x)\,\mathrm{d}\mu(x^{\prime}).}$
@@
We next show that, similar to classic OT, QOT satisfies stability with respect to the marginals.
@@
Stability is crucial in classic OT theory as it ensures that numerical algorithms for solving the OT problem converge consistently. On the other hand, numerically computing or approximating the QOT solution remains a difficult task, as the discretized version remains an NP-hard problem (Loiola et al. (2007)), a fact already noted in Mémoli (2011a, Remark 4.6). The recent work of Kravtsova (2024) also indicates the NP-hardness of the GW distance. As a discrete Monge version of QOT, QAP provides feasible heuristic algorithms; see Burkard et al. (2012) and Çela (2013). However, one needs to be careful here since discrete QOT may not always have a Monge minimizer even if a bijective transport map exists (see Example 1.1). We leave the computational aspects of QOT for further investigation.
@@
The lower bounds in Propositions 4.6 and 4.7 are generally not sharp, but they are useful in proving the optimality of some transport plans. The essential ideas of these lower bounds are present in the literature on QAP and GW distances. More precisely, Proposition 4.6 generalizes Lawler’s lower bound on QAP (Burkard et al. (2012)) and is known as the “third lower bound” in the GW setting (Mémoli (2007, 2011a)); Proposition 4.7 is an extension of the “second lower bound” for the GW distance. Examples 3.6 and 3.7 detail the connections between GW distance and QAP to QOT.